# Attention In Geometry: An Adaptive Density Field (ADF) with FAISS-Accelerated Spatial Attention

### v1

@@ Abstract
Spatial computation in geographic and urban systems increasingly demands scalable mechanisms for aggregating local influence under metric constraints. Classical approaches such as kernel density estimation [silverman1986density] and spatial interpolation either rely on global summation or treat approximation solely as an implementation concern, limiting their interpretability and scalability in large-scale settings. This work introduces Adaptive (Attention) Density Fields (ADF), a geometric attention framework that formulates spatial aggregation as a query-conditioned, metric-induced attention operator in continuous space. Given a set of spatial points with associated scalar scores, ADF constructs a local adaptive Gaussian mixture centered at the query’s nearest neighbors, where kernel bandwidths are modulated by point-specific scores. Approximate nearest-neighbor search, implemented via FAISS [johnson2017faiss] inverted file indices, is treated as an intrinsic component of the attention mechanism, enabling scalable execution while preserving locality. ADF bridges concepts from adaptive kernel methods and attention mechanisms by reinterpreting spatial influence as geometry-preserving attention, grounded directly in physical distance rather than learnt latent projections. The framework is formulation-level rather than algorithm-specific, allowing flexible kernel choices, score-to-bandwidth mappings, and approximation parameters. As such, ADF provides a unifying perspective on spatial influence modeling that emphasizes structure, scalability, and geometric interpretability, with relevance to geographic information systems, urban analysis, and spatial machine learning.
This manuscript is a preprint and has not undergone peer review.
1. Introduction
Large-scale spatial analysis increasingly involves aggregating millions of spatially distributed points of interest (POIs), each associated with heterogeneous influence scores. A common goal is to evaluate a continuous influence field $F(x)$ that reflects the accumulated effect of nearby points under physical distance constraints. However, naive kernel-based summation over all points is computationally infeasible at scale, and classical spatial interpolation methods often treat approximation purely as an implementation concern rather than a structural component of the model.
This work proposes a geometric attention framework for scalable continuous spatial aggregation. Instead of viewing spatial influence as a global density estimation problem, we formulate it as a query-conditioned, metric-induced attention operator in continuous space. Locality, sparsity, and approximation are treated as intrinsic to the formulation, rather than as post hoc accelerations.
This perspective is related in spirit to kernelized and geometry-aware attention mechanisms, but differs in that attention weights are induced directly by physical distance rather than learned projections. [bronstein2021geometric] [choromanski2020rethinking]
Given a set of POIs with associated scalar scores, the framework proceeds as follows:
  • Map POIs from geodetic $(\lambda,\varphi,h)$ to ECEF $\mathbf{x}\in\mathbb{R}^{3}$;
  • Associate each POI with a scalar score $s_{i}$ representing application-dependent influence;
  • Construct an inverted file (IVF) index to accelerate approximate $k$-nearest-neighbor (k-NN) search in 3D;
  • For any query location $\mathbf{x}$:
    • retrieve its $k$ nearest neighbors,
    • define a Gaussian kernel centered at each neighbor, with bandwidth modulated by its score,
    • aggregate the weighted kernel contributions.
The resulting field $F(\mathbf{x})$ constitutes an adaptive Gaussian density/intensity field (ADF): a finite, query-dependent Gaussian mixture that enables scalable evaluation of continuous spatial influence over millions of points with sublinear query complexity.
Importantly, this work focuses on the formulation of geometric attention for spatial aggregation rather than any single algorithmic instantiation. The choice of kernel family, score-to-bandwidth mapping, normalization, and approximation parameters are treated as design degrees of freedom. The emphasis is on the structural decomposition of spatial attention—query-conditioned neighbor selection, metric-grounded weighting, and efficient execution—providing a unifying perspective on adaptive kernel methods, spatial indexing, and attention mechanisms in continuous space.
2. Framework
@@
2.1. Overview
The Adaptive Density Field (ADF) is a geometric attention operator defined over continuous three-dimentional space. Given a set of spatial points of interest (POIs) with associated scalar scores, ADF constructs a continuous influence field by aggregating local contributions from nearby points.
For each query location, ADF performs a query-conditioned neighbor selection using an approximate k-NN search, followed by a score-modulated kernel weighting in metric space. The aggregation is explicitly localized, sparse, and adaptive, with kernel bandwidths determined by POI scores rather than fixed global parameters.
To ensure scalability to larger datasets, ADF integrates approximate nearest-neighbor search via FAISS inverted file indices, reducing per query complexity from linear to sublinear in the number of POIs. The resulting operator produces a continuous scalar field that can be evaluated at arbitrary spatial locations and serves as a geometry-preserving attention mechanism grounded in physical distance.
2.2. Data Preparation
Suggested data structures: positions and scores
In the experiments, POIs are derived from a physics-informed trajectory analysis pipeline that detects anomalous motion based on time-normalized Mahalanobis [mahalanobis1936distance] residuals. Each POI is associated with a spatial location and a scalar score reflecting deviation magnitude. The details of this pipeline are not central to the ADF formulation and are therefore omitted; ADF operates on arbitrary spatial point sets with associated scores, defined as follows:
  • $\mathbf{x}_{i}\in\mathbb{R}^{3}$: ECEF position of POI $i$
  • $s_{i}$: score (weight) of POI $i$ Formally:
    $\{\mathbf{x}_{i},s_{i}\}_{i=1}^{n}$
    with:
    $\mathbf{x}_{i}\in\mathbb{R}^{3},\quad s_{i}\in\mathbb{R}.$
    The score $s_{i}$ represents an application-dependent measure of influence, saliency, or importance, and its semantic interpretation is not constrained by the ADF formulation.
2.3. Methodolody
With the POI data, we can perform the ADF construction with the following method:
2.3.1. FAISS IVF Index: Accelerating Neighbor Search
For each POI data we imported, we assign the IVF index [jegou2011product] , mathematically:
Training:
  • FAISS runs k-means with $n_{\text{list}}=4096$ clusters in $\mathbb{R}^{3}$.
  • It learns centroids $\{\mathbf{c}_{j}\}_{j=1}^{4096}$.
    $\min_{\{\mathbf{c}_{j}\}}\sum_{i=1}^{n}\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$
Assignment: Each vector $\mathbf{x}_{i}$ is assigned to its nearest centroid:
$\ell(i)=\arg\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$
@@ 2.3.1. FAISS IVF Index: Accelerating Neighbor Search
Search: For a query $\mathbf{x}$, FAISS finds the 16 nearest centroids:
$j_{1},\dots,j_{16}=\text{indices of the 16 closest }\mathbf{c}_{j}\text{ to }\mathbf{x}.$
and Searches only in the union of lists:
$L_{j_{1}}\cup\dots\cup L_{j_{16}}$
2.3.2. ADF model: an adaptive gaussian mixture
@@
For a query point $\mathbf{x}\in\mathbb{R}^{3}$ FAISS returns indices:
$\{i_{1},\dots,i_{k}\}=\text{k-NN}_{\text{IVF}}(\mathbf{x})$
in approximate nearest‑neighbor sense. Then set:
  • Neighbor positions:
    $\mathbf{x}_{i_{j}}\in\mathbb{R}^{3}$
  • Neighbor scores:
    $s_{i_{j}}$
@@
Mathematically, for each neighbor $j=1\dots k$:
$\mathbf{d}_{j}=\mathbf{x}_{i_{j}}-\mathbf{x}=\begin{bmatrix}x_{i_{j},1}-x_{1}\\ x_{i_{j},2}-x_{2}\\ x_{i_{j},3}-x_{3}\end{bmatrix}\in\mathbb{R}^{3}.$
@@
For each neighbor $j$:
$\sigma_{j}=\frac{\sigma_{0}}{s_{i_{j}}+10^{-6}}.$
@@
  • Low score $s_{i_{j}}\implies$ larger $\sigma_{j}\implies$ broader kernel. This is why it’s adaptive: kernel width depends on the score. Then for each $j$ we have:
    $\text{inv\_var}_{j}=\frac{1}{\sigma_{j}^{2}}.$
NOTE: The specific functional form relating scores to bandwidths (Eq. 5) is a design choice, not a defining property of ADF. ADF is defined by (i) query-conditioned neighbor selection, (ii) metric-induced kernel weighting, and (iii) scalable approximate execution. Any monotonic or learned score-to-bandwidth mapping may be substituted without altering the underlying framework.This work focuses on the general framework of adaptive, score-modulated kernels; alternative monotonic mappings or learned parameterizations can be substituted without altering the structure of ADF.
Quadratic Form (Mahalanobis Distance Squared)
From equation 4 (Eq. 4) we can derive the componentwise square:
$(\mathbf{d}_{j})^{2}=\begin{bmatrix}(x_{i_{j},1}-x_{1})^{2}\\ (x_{i_{j},2}-x_{2})^{2}\\ (x_{i_{j},3}-x_{3})^{2}\end{bmatrix}$
Additionally, we have an array with shape $(k,1)$, broadcasts $\text{inv\_var}_{j}$ across the 3 dimensions. So for each neighbor $j$:
$\text{quad}_{j}=\sum_{d=1}^{3}(x_{i_{j},d}-x_{d})^{2}\cdot\frac{1}{\sigma_{j}^{2}}=\frac{\|\mathbf{x}_{i_{j}}-\mathbf{x}\|^{2}}{\sigma_{j}^{2}}.$
This is exactly the Mahalanobis distance squared for an isotropic Gaussian with variance $\sigma_{j}^{2}$:
$\text{quad}_{j}=(\mathbf{x}-\mathbf{x}_{i_{j}})^{T}\Sigma_{j}^{-1}(\mathbf{x}-\mathbf{x}_{i_{j}}),$
@@
$K_{j}(\mathbf{x})=\exp\left(-\frac{1}{2}\text{quad}_{j}\right)=\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$
In this work, we assume isotropic kernels with $\Sigma_{j}=\sigma_{j}^{2}I_{3}$. And anisotropic kernels are discussed as a future extension. Then the ADF value at $\mathbf{x}$ is:
$F(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\,K_{j}(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$
Therefore, we claim that ADF is a finite Gaussian mixture [bishop2006pattern] centered at the nearest neighbors, with:
  • centers $\mathbf{x}_{i_{j}}$
  • weights $s_{i_{j}}$
@@
Assume (i) $s_{i}\geq 0$, or $s_{i}$ is transformed via $s_{i}\leftarrow\text{softplus}(s_{i})$, (ii) $\sigma_{0}$ or the score-to-bandwidth mapping can be learned via gradient descent, (iii) $\sigma_{0}$ has units of meters and controls the physical spatial influence scale.
This can be interpreted as an adaptive kernel density / intensity field in 3D ECEF space.
The contribution of this work lies in the formulation of ADF as a geometric attention mechanism with scalable execution. Kernel choices, bandwidth parameterizations, and score transformations are intentionally left flexible to emphasize the generality of the framework.
2.4. Computational complexity analysis
2.4.1. Time complexity
Without FAISS:
$O(n)\text{ per query}$
With FAISS IVF:
$O(n_{\text{probe}}\cdot\frac{n}{n_{\text{list}}}+k)$
2.4.2. Memory Complexity
Memory complexity is
$O(n)$
dominated by the inverted lists.
2.5. Evaluation
2.5.1. Approximation Error Acknowledgement
Since FAISS IVF performs approximate nearest-neighbor search, ADF(x) is an approximation of the full kernel sum. However, locality of the Gaussian kernel ensures distant errors contribute negligibly. In practice, approximation quality is controlled by $n_{\text{probe}}$ and $k$.
2.5.2. Evaluating ADF at all POIs
For each POI position $\mathbf{x}_{i}$:
$\text{ADF}_{i}=F(\mathbf{x}_{i})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}_{i}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right),$
where $\{i_{1},\dots,i_{k}\}$ are the $k$-nearest neighbors of $\mathbf{x}_{i}$ according to the IVF‑accelerated FAISS search. So the final result is a scalar ADF field defined on all your POIs:

### v2

@@ Abstract
Spatial computation in geographic and urban systems increasingly requires query-conditioned, local, interpretable aggregation under metric constraints. Many classical approaches and spatial interpolations rely on global summation and treat approximation primarily as an implementation concern, which limits their interpretability and scalability in large-scale settings. We put forward Adaptive Density Field (ADF) as a geometric attention framework that formulates spatial aggregation as a query-conditioned, metric-induced attention operator in continuous space. Given a set of labelled spatial points with associated scalar scores, ADF defines a continuous influence/intensity field over space. For a given query location, the field value is obtained via a local adaptive Gaussian kernel mixture centered on the query’s nearest neighbors, where kernel bandwidths are modulated by point-specific scores, such that we can evaluate a scalar value representing local aggregated influence. Additionally, approximate nearest-neighbor search is introduced as a complement, which can be implemented with various acceleration algorithms and enables scalable execution while preserving locality. The proposed ADF bridges concepts from adaptive kernel methods, classical GIS methods, and the attention mechanisms by reinterpreting spatial influence as geometry-embedded attention, grounded directly in physical distance rather than learnt latent projections. The proposed framework is formulation-level rather than algorithm-specific, allowing flexible kernel choices, score-to-bandwidth mappings and approximation parameters. Thus, this approach provides a unifying perspective on spatial influence modeling that emphasizes structure, scalability, and geometric interpretability, with relevance to geographic information systems, urban analysis, and spatial machine learning.
1 Introduction
Geographic Information Science (GIScience) is a multidisciplinary field [14] concerned with developing analytical and predictive methods for specific tasks involved with geographic data. While recent advances in computer science have been dominated by the thriving of large language models [6], people often ignore that a large-scale spatial systems also require a query-conditioned, interpretable aggregation method, like the ones proposed for large language models, that respects metric locality while remaining computationally feasible. Such requirements arise naturally in a range of GIScience scenarios, including on-demand accessibility estimation in urban environments, real-time assessment of spatial influence around moving objects (e.g., vehicles, drones, or trajectories), point-of-interest impact modeling, adaptive exposure analysis in environmental monitoring, and interactive spatial querying in large-scale geographic databases [33, 21, 23, 29]. In these settings, spatial influence must be evaluated locally at arbitrary query locations, often under strict latency and memory constraints, which makes global or batch-based aggregation methods impractical [28].
Classical adaptive KDE [30] and k-NN density estimator [10] provide a statistical foundation for modern variable bandwidth smoothing, but they are typically global or batch estimators and do not treat approximation as part of the operator definition [4]. Currently, on one hand several recent applied pipelines have attempted to combine ANN acceleration with KDE for large datasets for better scalability [32, 19, 38], and on the other hand, query-driven KDE variants have also been proposed for influence region computation and other domains [2, 15]. However, these approaches either assume full density estimation or lack explicit score-modulated attention semantics and intrinsic sparsification, which limits their ability to explicitly encode query-conditioned sparsification, score-dependent influence modulation, and operator-level interpretability. In particular, approximation is typically treated as an external optimization rather than an intrinsic component of the aggregation operator, and kernel influence is not formulated as an attention-like mechanism that directly reflects spatial locality and point-specific importance. As a result, while being robust and well-established, existing methods unavoidably offer limited flexibility in trading off locality, interpretability, and computational efficiency within a unified formulation.
In this paper, we propose the Adaptive Density Field (ADF), a geometry‑embedded attention operator that makes locality, sparsification, and approximation first‑class design choices. In particular, we consider the problem of constructing a continuous influence field $F(x)$ that reflects the accumulated effect of nearby points of interest (POIs) under physical distance constraints. This framework formulates spatial aggregation as a query-conditioned, metric-driven weighted sum. For any query location $x\in\mathbb{R}^{3}$ (Earth-Centered, Earth-Fixed coordinates), ADF selects a local neighbor set, assigns each neighbor a score-modulated kernel, and aggregates the contributions to produce $F(x)$, using the FAISS inverted file indices [18] for acceleration. This perspective is related to kernelized and geometry-aware attention mechanisms, but differs in that attention weights are induced directly by physical distance rather than learned projections [7, 9].
This work is structured as follows: In Section 2, we introduce the basic methodology and formal definition of the Adaptive Density Field (ADF), positioning the framework as a conceptual contribution to GIScience. In Section 3, we present a practical instantiation that combines ADF with ANN indices for POI detection. In Section 4 we evaluate the robustness and efficiency of the proposed framework, and demonstrate the selection of parameters. In Section 5 we discuss the relationship and differences between the proposed method and several classic and well-known GIS method. In Section 6, we discuss the implications of ADF as a unifying operator-level perspective on spatial influence modeling, emphasizing geometric grounding, explicit approximation, and scalability. Finally, Section 7 summarizes the main conclusions and outlines future work.
2 Framework
2.1 Overview
The Adaptive Density Field (ADF) is a geometric attention operator designed to unify locality, sparsity, and approximation in continuous spatial aggregation. Rather than focusing on implementation or dataset-specific optimizations, the framework formalizes spatial influence as an operator-level construct: given a query location $x\in\mathbb{R}^{3}$ (ECEF coordinates), ADF aggregates contributions from nearby points of interests (POIs) through score-modulated kernels and a metric-induced weighting scheme.
This section provides a conceptual roadmap of the framework. We first define the operator mathematically, discuss the score-to-bandwidth mappings, and approximation mechanisms, and then present the usage of FAISS for efficient neighbor retrieval as an approximate nearest-neighbor (ANN) method. Finally, we evaluate the time and memory complexity of the proposed method, highlighting the flexibility of the approach and its suitability for a range of GIScience applications.
2.2 Data Preparation
Suggested data structures: positions and scores
In the experiments, POIs are derived from a physics-informed trajectory analysis pipeline (detailed in the Appendix); each POI is associated with a spatial location and a scalar score reflecting deviation magnitude. The details of this pipeline are not central to the ADF formulation and are therefore omitted. ADF operates on arbitrary spatial point sets with associated scores, defined as follows:
(i) $\mathbf{x}_{i}\in\mathbb{R}^{3}$ is the ECEF position of POI $i$
(ii) $s_{i}$ is the score (weight) of POI $i$. Formally, the input is defined as $\{\mathbf{x}_{i},s_{i}\}_{i=1}^{n}$, where $\mathbf{x}_{i}\in\mathbb{R}^{3}$ and $s_{i}\in\mathbb{R}$. The score $s_{i}$ represents an application-dependent measure of influence, saliency, or importance; its semantic interpretation is not constrained by the ADF formulation.
@@ 2.2 Data Preparation
2.3 Methodology
Given the POI data, we construct the ADF as follows:
2.3.1 FAISS IVF Index: Accelerating Neighbor Search
For the proposed framework, FAISS is used purely as a computational accelerator to retrieve local neighborhoods efficiently from large POI sets. Parameters such as k and nprobe control access to sufficient local context rather than defining the ADF model itself (we will discuss the parameter selection in Section 4). In practice, results were stable once neighborhoods exceeded a minimal size, and all experiments used a fixed, conservative configuration. For each POI data imported, we assign the IVF index [17], and the procedure for high-dimensional indexing is formalized as follows:
Training
(i) FAISS runs k-means with $n_{\text{list}}=4096$ clusters in $\mathbb{R}^{3}$.
(ii) It learns centroids $\{\mathbf{c}_{j}\}_{j=1}^{4096}$.
This formalizes the cluster assignment, though standard FAISS documentation may be consulted for full details.
Assignment
Each vector $\mathbf{x}_{i}$ is assigned to its nearest centroid:
$\ell(i)=\arg\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$ (1)
@@ 2.3.1 FAISS IVF Index: Accelerating Neighbor Search
Search
For a query $\mathbf{x}$, FAISS finds the 16 nearest centroids $j_{1},\dots,j_{16}$: indices of the 16 closest $\mathbf{c}_{j}$ to $\mathbf{x}.$ Then searches only in the union of these lists: $L_{j_{1}}\cup\dots\cup L_{j_{16}}$.
2.3.2 ADF Model: An Adaptive Gaussian Mixture
@@
For a query point $\mathbf{x}\in\mathbb{R}^{3}$ FAISS returns indices: $\{i_{1},\dots,i_{k}\}=\text{k-NN}_{\text{IVF}}(\mathbf{x})$ in approximate nearest‑neighbor sense. Then set: (i) Neighbor positions: $\mathbf{x}_{i_{j}}\in\mathbb{R}^{3}$; (ii) Neighbor scores: $s_{i_{j}}$
@@
Mathematically, for each neighbor $j=1\dots k$, we have:
$\mathbf{d}_{j}=\mathbf{x}_{i_{j}}-\mathbf{x}\in\mathbb{R}^{3}.$ (2)
@@
Conceptually, higher score points are supposed to represent stronger local influence, hence narrower kernels; lower scores should spread influence more diffusely. Therefore, for each neighbor $j$:
$\sigma_{j}=\frac{\sigma_{0}}{s_{i_{j}}+10^{-6}}.$ (3)
@@
  • Low score $s_{i_{j}}\implies$ larger $\sigma_{j}\implies$ broader kernel. This is why it’s adaptive: kernel width depends on the score.
Then for each $j$ we have:
$\text{inv\_var}_{j}=\frac{1}{\sigma_{j}^{2}}.$ (4)
NOTE: The specific functional form relating scores to bandwidths is a design choice, not a defining property of ADF. ADF is defined by (i) query-conditioned neighbor selection; (ii) metric-induced kernel weighting; and (iii) scalable approximate execution. While this study employs a specific reciprocal mapping, any monotonic or learned parameterization may be substituted without altering the underlying framework.
Quadratic form (Mahalanobis distance squared [22])
From Equation 3 we can derive the componentwise square $\mathbf{d}_{j}^{2}$. Additionally, we have an array with shape $(k,1)$, broadcasts $\text{inv\_var}_{j}$ across the 3 dimensions. So for each neighbor $j$:
$\text{quad}_{j}=\sum_{d=1}^{3}(x_{i_{j},d}-x_{d})^{2}\cdot\frac{1}{\sigma_{j}^{2}}=\frac{\|\mathbf{x}_{i_{j}}-\mathbf{x}\|^{2}}{\sigma_{j}^{2}}.$ (5)
This is a derived Mahalanobis distance squared for an isotropic Gaussian with variance $\sigma_{j}^{2}$:
$\text{quad}_{j}=(\mathbf{x}-\mathbf{x}_{i_{j}})^{T}\Sigma_{j}^{-1}(\mathbf{x}-\mathbf{x}_{i_{j}}),$ (6)
@@
$K_{j}(\mathbf{x})=\exp\left(-\frac{1}{2}\text{quad}_{j}\right)=\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$ (7)
In this work, we assume isotropic kernels with $\Sigma_{j}=\sigma_{j}^{2}I_{3}$. While the ADF framework natively supports anisotropic covariance matrices to account for directional trajectory influence, for simplification, we applied isotropic kernels in this study to establish a controlled baseline for evaluating the scalability and partitioning stability of the ADF. This isotropic simplification isolates the effects of score-modulated adaptation and scalable neighbor retrieval, establishing a robust baseline before introducing directional anisotropy in future trajectory-specific extensions. Then the ADF value at $\mathbf{x}$ is assumed:
$F(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\,K_{j}(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$ (8)
Note that $F(x)$ is not normalized to integrate to unity and is therefore not a probability density.
Therefore, we claim that ADF is a finite Gaussian mixture [5]. Here, the term ‘Gaussian mixture’ refers to a computational representation rather than a probabilistic generative model, centered at the nearest neighbors, with:
  • centers $\mathbf{x}_{i_{j}}$;
  • weights $s_{i_{j}}$;
@@
We assume (i) $s_{i}\geq 0$, or $s_{i}$ is transformed via $s_{i}\leftarrow\text{softplus}(s_{i})$, (ii) $\sigma_{0}$ or the score-to-bandwidth mapping can be learned via gradient descent, (iii) $\sigma_{0}$ has units of meters and controls the physical spatial influence scale. This can be interpreted as an adaptive kernel influence / intensity field in 3D ECEF space.
The contribution of this work lies in the formulation of ADF as a geometric attention mechanism with scalable execution. Kernel choices, bandwidth parameterizations, and score transformations are intentionally left flexible to emphasize the generality of the framework. In future extensions tailored to directional flow data (e.g., aircraft trajectories), anisotropic covariances can be induced by aligning principal axes with local velocity vectors, fully realizing vector-aware geometric attention absent in traditional VBKDE.
2.3.3 Evaluating ADF at all POIs
From Equation 8, for each POI position $\mathbf{x}_{i}$:
$\text{ADF}_{i}=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}_{i}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right),$ (9)
where $F(x)$ evaluated at POI $x_{i}$ is denoted as $ADF_{i}$, and $\{i_{1},\dots,i_{k}\}$ are the $k$-nearest neighbors of $\mathbf{x}_{i}$ according to the IVF‑accelerated FAISS search. So the final result is a scalar ADF field defined on all your POIs: